OpenStudy (anonymous):

Use an addition property to solve for c. 14 + (–12) = –12 + c A. –26 B. –14 C. 14 D. 26

OpenStudy (anonymous):

will give metals!

OpenStudy (anonymous):

c=14

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

What value of k makes the statement true? –(k + 12) = 6 + (–12) A. –18 B. –6 C. 6 D. 18

OpenStudy (anonymous):

can you do this^

OpenStudy (anonymous):

k=-6

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

this to please: What is the value of the expression? 12.4 + (–2 + 1.6) A. 11 B. 12 C. 15 D. 16

OpenStudy (anonymous):

12

OpenStudy (anonymous):

u welcoe

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

this>>>>>>> Which is an example of the associative property? A. –(18 + 3) = –18 + (–3) B. –17 + 0 = –17 C. (–8 + 10) + 4. = –8 + (10 + 4) D. –5 + 7 = 7 + (–5)

OpenStudy (anonymous):

@famasow

OpenStudy (anonymous):

c